OpenStudy (anonymous):

What is 3 √250 expressed in simplest radical form?

OpenStudy (lgbasallote):

\[\sqrt{250} = \sqrt{25 \times 10} \longleftarrow \text{25 is a perfect square}\]

OpenStudy (unklerhaukus):

\[3 √250=3\times\sqrt{25\times10}\]\[=3\times\sqrt{5^2\times2\times5}\]

OpenStudy (anonymous):

soo is the answer 5/10

OpenStudy (unklerhaukus):

nope

OpenStudy (unklerhaukus):

\[=3×\sqrt{5^2×2×5}=3\times5\times\sqrt 2\times \sqrt 5\]

OpenStudy (anonymous):

but 5 times 5 is 25

OpenStudy (unklerhaukus):

dont forget the three danesmart

OpenStudy (anonymous):

how i get the 15

OpenStudy (lgbasallote):

\[3\sqrt{250} = 3\sqrt{25 \times 10} = 3 \times \sqrt{25} \times \sqrt{10}\] 25 is a perfect square so you can put it out \[3 \times 5 \times \sqrt 10\] got it??

OpenStudy (anonymous):

oooooooooooooooooooo ok thank you now I get it :D

OpenStudy (unklerhaukus):

the answer is fifteen times the square root of two times the square root of 5